hartnn (hartnn):

Find the stationary values of sinx sin y sin (x+y). To get stationary values, we solve f_x=0 and f_y=0 simultaneously. f_x=0= cos x sin y sin (x+y)+sin x sin y cos (x+y) f_y=0= sin x cos y sin (x+y)+sin x sin y cos (x+y) Adding the above 2 equations, sin (x+y) [cos x sin y+sin x cos y ]+cos (x+y) sin x sin y=0 sin^2 (x+y)+cos (x+y) sin x sin y=0 …. what next ? or any other better approach ?

hartnn (hartnn):

@SithsAndGiggles

OpenStudy (freckles):

\[\cos(x)\sin(y)\sin(x+y)=\sin(x)\cos(y)\sin(x+y) \\ \sin(x+y)(\sin(x)\cos(y)-\sin(y)\cos(x))=0 \\ \sin(x+y)\sin(x-y)=0\]

OpenStudy (freckles):

set both factors =0

OpenStudy (freckles):

sin(x+y)=0 or sin(x-y)=0

OpenStudy (freckles):

@hartnn any questions? I think you taught to hard.

OpenStudy (freckles):

too*

hartnn (hartnn):

i got that, but there will be other values too , right ? i need all the stationary values. from that i get x+y =0, x-y =0 hence x=y=0

OpenStudy (freckles):

I get x+y=npi and x-y=npi therefore x=npi and y=0

OpenStudy (freckles):

\[2x+0=2n \pi => x=n \pi => y=0 \]

OpenStudy (freckles):

sin(x+y)=0 sin(u)=0 when u=npi so x+y=npi sin(x-y)=0 sin(v)=0 when v=npi so x-y=npi _______________________add equations 2x=2npi

hartnn (hartnn):

-2pi/3 , 2pi/3 -2pi/3 , pi/3 and all other values....how would i get them ?

OpenStudy (freckles):

hmm...

OpenStudy (freckles):

-x-y=npi x-y=npi --------- -2y=2npi and I guess we can also get y=-npi when x=0 but for the other values thinking

OpenStudy (freckles):

\[0=\sin(x+y)(\sin(x)\cos(y)+\sin(y)\cos(x)) \\+\cos(x+y)(\sin(x)\sin(y)+\sin(x)\sin(y)) \\ 0=\sin(x+y)\sin(x+y)+\cos(x+y)2\sin(x)\sin(y)\]

OpenStudy (freckles):

i notice in your equation above you don't have that 2 constant multiple for the second addend

hartnn (hartnn):

oh yeah, missed it!

hartnn (hartnn):

sin^2 (x+y)+2 cos (x+y) sin x sin y=0

OpenStudy (freckles):

\[0=1-\cos^2(x+y)+\cos(x+y)2\sin(x)\sin(y) \\ \cos^2(x+y)-\cos(x+y)2\sin(x)\sin(y)-1=0\] it would be easier if that 2sin(x)sin(y) wasn't there :(

hartnn (hartnn):

yeah, a quadratic in cos (x+y) but that would still have given values of x+y

OpenStudy (freckles):

\[\sin(x)\sin(y)=\frac{1-\cos^2(x+y)}{-2\cos(x+y)}=\frac{-\sin^2(x+y)}{2\cos(x+y)}\] \[\frac{-\sin^2(x+y)}{2\cos(x+y)} \sin(x+y)=0\] This doesn't work either ... I tried solving our equation for sin(x)sin(y) and pluggin it into the original where we sit the original=0 but this still only gives us sin(x+y)=0

OpenStudy (freckles):

\[\cos(x+y)=\frac{-2\sin(x)\sin(y) \pm \sqrt{4\sin^2(x)\sin^2(y)+4}}{2}\] so \[\cos(x+y)=\frac{-2\sin(x)\sin(y) \pm 2 \sqrt{\sin^2(x)\sin^2(y)+1}}{2}\] \[\cos(x+y)=-\sin(x)\sin(y) \pm \sqrt{(\sin(x)\sin(y))^2+1}\] I don't think I know what to do with this if i can do anything with it just thinking

OpenStudy (anonymous):

I wasn't sure if anyone brought it up, but we could use an identity for the \(f_x=0\) equation: \[\begin{align*} f_x&=\cos x\sin y\sin (x+y)+\sin x\sin y\cos(x+y)\\ 0&=\cos x\sin (x+y)+\sin x \cos (x+y)\\ 0&=\sin(2x+y)&\text{(angle sum id.)} \end{align*}\] Similarly for \(f_y=0\), \[\begin{align*} f_y&=\sin x\cos y\sin (x+y)+\sin x\sin y\cos(x+y)\\ 0&=\cos y\sin (x+y)+\sin y\cos(x+y)\\ 0&=\sin(x+2y)&\text{(angle sum id.)} \end{align*}\] So we have \[\begin{cases} 2x+y=n\pi\\ x+2y=n\pi \end{cases}~~\implies~~x=y\] hmm, does that help with anything?

OpenStudy (freckles):

yeah he already has the x=y he wants to other solutions wolfram gave

OpenStudy (freckles):

but i like the way you found that solution it is kinda cuter than what i did

OpenStudy (freckles):

well I think 2x+y=npi and x+2y=npi actually gives us your other solutions @hartnn 2x+y=npi -2x-4y=-2npi ------------------------- -3y=-2npi y=2npi/3 then x=npi-2(2npi/3)=-npi/3

hartnn (hartnn):

awesome! thank you very much :)